Water entering a house flows with a speed of 0.25 m/s through a pipe of 1.1 cm inside radius. What is the speed of the water at a point where the pipe tapers to a radius of 2.3 mm?

__________m/s
Answer #1

here ,

let the speed of particle at the taper end is v2

using equation of continuity

v1 * A1 = v2 * A2

0.25 * pi*(1.1)^2 = v2 * pi*(.23)^2

v2 = 5.71 m/s

speed of the water at a point where the pipe tapers to a radius of 2.3 mm is 5.71 m/s